# Jacobi’s solution for geodesics on a triaxial ellipsoid

### v1

@@ 1 Introduction
The geodesic is the shortest path between two points on a surface, and it plays a crucial role in geodesy, where the earth is typically modeled as an ellipsoid of revolution, a biaxial ellipsoid. The main geodesic problems are (1) given a starting point and a direction, find the point a certain distance away, the “direct” geodesic problem, and (2) to find the length and direction of the geodesic connecting two points, the “inverse” geodesic problem. The path of a geodesic is also given by the motion of a point mass sliding on the surface of the ellipsoid without friction and in the absence of external forces. This allows geodesics to be extended indefinitely.
Recently, there has been interest in using a refined model of the earth as a triaxial ellipsoid (Panou et al., 2020). Consequently, there is interest in obtaining reliable solutions to the geodesic problem in this case.
@@
The triaxial ellipsoid, on the other hand, possesses no obvious symmetry. It therefore came as a surprise when Jacobi (1839) found that the geodesic problem could be reduced to quadrature in this case too, with the solution given in terms of one-dimensional integrals. (We know the date, given in the abstract, for this discovery, because of a letter he wrote on December 28, 1838, to F. W. Bessel, his neighbor in Königsberg.) Jacobi (1843, §28) expanded on his method in his Lectures on Dynamics, and the result was generalized by Liouville (1846, §§20–21) to apply to so-called Liouville surfaces. The qualitative properties of the solution can be found in several textbooks (Darboux, 1894; Hilbert and Cohn-Vossen, 1952; Klingenberg, 1982; Arnol’d, 1989; Berger, 2010). However, unlike the case of the biaxial ellipsoid, where Bessel (1825) provided a prescription for computing geodesics, little effort was given to implementing Jacobi’s solution.
@@ 2 Ellipsoidal coordinates
$\mathbf{V}=\sin\alpha\hat{\mathbf{E}}+\cos\alpha\hat{\mathbf{N}}$, (7)
where $\alpha$ is the azimuth of the geodesic measured clockwise from a line of constant $\omega$. At the pole of an oblate ellipsoid, we take the limit $\cos\beta\rightarrow 0+$, to give
@@
[…]
@@ 3 Qualitative behavior
For a biaxial ellipsoid, the equator and all the meridians are simple (not self-intersecting) closed geodesics. On the other hand, for a triaxial ellipsoid (provided it is not too eccentric), there are only three simple closed geodesics, namely the 3 principal ellipses. The major and minor ellipses are stable; if they are perturbed, the resulting geodesic oscillates about the original ellipse. However, the median ellipse is unstable; if the geodesic is perturbed, it swings away from the $Y=0$ plane before returning to the original ellipse, but now traveling along it in the opposite direction. The stability of closed geodesics is treated in Appendix F.
@@ 4 Jacobi’s solution
[…]
[scale=0.75]figures/cayley
[…]
[scale=0.75]figures/cayleydist
@@
$\frac{s+s_{1}}{b}$ $=\int\frac{k\cos^{2}\beta\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
$\quad{}\pm\int\frac{k^{\prime}\sin^{2}\omega\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega.$ (17b)
@@ 5 Properties of general geodesics
The course of a geodesic is determined by Eq. (19a) and we are immediately confronted with the problem that, for $\gamma\neq 0$, one of the integrands in Eq. (17) is singular and that, if $\gamma=0$, the integrals themselves are singular. In this section, we address the circum- and transpolar cases where $\gamma\neq 0$.
@@ 7 Biaxial ellipsoids
For biaxial ellipsoids, the functions $f_{\theta}$ and $g_{\theta}$ are evaluated with $\kappa=0$ to give
@@
The functions $f_{\psi}$ and $g_{\psi}$, with $\kappa=1$, become
@@
In the spherical limit, we carry out the integration to give
@@ 10 The inverse problem
[…]
[scale=0.75]figures/cuta
[…]
[scale=0.75]figures/cutb
@@
[…]
[scale=0.75]figures/cutunroll
[…]
[scale=0.75]figures/hybrid
@@
  • With the biaxial problem, Helmert (1880, Eq. (6.5.1)) provided a formula for the reduced length $m_{12}$, allowing conjugate points to be found by the condition, $m_{12}=0$. I know of no corresponding formula for $m_{12}$ for the triaxial case. Nevertheless, it is possible to find the first conjugate point by the condition that $\beta$ has completed half an oscillation, which in turn implies that $\psi$ has advanced by $\pi$. Equation (26a) can be used to give the value of $\omega$ at that point. In the special case of a biaxial ellipsoid, the conjugate point for a meridional geodesic is given by Eq. (49); this is only needed for solving inverse problems on a prolate ellipsoid.
@@
11 Implementation
An implementation of the solutions of the direct and inverse geodesic problems is provided with version 2.6 of the C++ library, GeographicLib (Karney, 2025). It’s practically impossible to exhaustively document the methods here; the reader is referred to the code for details. Here, I give an overview of the important aspects of the code.
@@ 11 Implementation
Testing for geodesics on a triaxial ellipsoid is about two orders of magnitude more challenging than the biaxial case. The shape of the ellipsoid is specified by 2 parameters ($e$ and $k$) instead of just one (the flattening), and the solution of the inverse problem depends on the longitudes of both endpoints instead of just their difference. So I limited my initial testing to a single ellipsoid, Cayley’s ellipsoid scaled to the median semiaxis, i.e., $a,b,c]=[\sqrt{2},1,1/\sqrt{2}]$, with an emphasis on exploring all the different inverse problems outlined in Sec. 10. The test set (Karney, 2024c) contains 500 000500\,000 geodesics, computed at high precision, with the coordinates of the endpoints given as integer degrees. My testing also included other ellipsoids with $a/c=2$, including oblate and prolate ellipsoids. In addition, I tested with spheres ($a=c$) with different values of $k$, and with a triaxial model of the earth (Panou et al., 2020). In the following, I will only report results for the published dataset for Cayley’s ellipsoid. These should be regarded as rough indications of those likely to be obtained with ellipsoids which are not too eccentric.
@@ Appendix C Newton with bisection in two dimensions
It turns out that for this class of two-dimensional root-finding problems, it is quite easy to establish an axis-aligned bounding rectangle, $x\in[x_{a},x_{b}]$ and $y\in[y_{a},y_{b}]$. This starts by including all of the initial bounding parallelogram—a much looser bound. On each iteration, update one edge of the bounding rectangle depending on the signs of $f$ and $g$. Thus, in the example shown in Fig. 7, the corners of the initial bounding box are marked by circles, and the initial guess, marked by a cross, is at its center. At this point, we find that $f<0$ and $g>0$. From the figure, it is clear that the desired solution lies below this point, and therefore, we update $y_{b}$, giving a new bounding box shown as the light rectangle. This is possible because of the constraints on the slopes of the curves $f=0$ and $g=0$, which, in turn, result from the form of these functions in Eqs. (C.1). With other sign combinations for $f$ and $g$, the other edges of the bounding rectangle can be updated, and in the case where either $f$ or $g$ vanishes, two edges can be updated.
@@ Appendix E Ordinary differential equations for geodesics
The equation for geodesics on a surface is the same as for the motion of a particle constrained to move on the surface but subject to no other forces. The centrifugal acceleration of the particle is $(V^{2}/R_{c})\hat{\mathbf{U}}$, where $R_{c}$ is the radius of curvature in the direction of the velocity $\mathbf{V}$. We will take the speed to be unity (and, of course, the speed is a constant in this problem); thus, time can be replaced by $s$, the displacement along the geodesic, as the independent variable. The differential equations for the geodesic are
@@
These errors in the solution given by integrating the ordinary differential equations are an impediment to solving the inverse problem. This depends sensitively on certain properties of the solution, e.g., that a geodesic leaving an umbilic intersects the opposite umbilic. I have stitched up a solution for the inverse problem in the Octave/MATLAB version of GeographicLib by tracking the solution in ellipsoidal coordinates as the solution in cartesian coordinates unfolds (Karney, 2024b). But the code is somewhat messy, because it involves repeated coordinate conversions and has to work around the small, but inevitable, errors in the solution of the direct problem.
@@ Appendix F The stability of closed geodesics
Test data is available at Karney (2024c). Additional test data is available upon request.
@@
A C++ implementation of the algorithms given in this paper is given in GeographicLib, version 2.6, available at https://github.com/geographiclib/geographiclib/releases/tag/r2.6. Version 2.7 of GeographicLib will be released soon; this will fix bugs found during the preparation of this manuscript.

### v2

@@ 1 Introduction
The shortest path between two points on a surface is a geodesic, and it plays a crucial role in geodesy, where the earth is typically modeled as an ellipsoid of revolution, a biaxial ellipsoid. The main geodesic problems are (1) given a starting point and a direction, find the point a certain distance away, the “direct” geodesic problem, and (2) to find the length and direction of the geodesic connecting two points, the “inverse” geodesic problem. The path of a geodesic is also given by a point mass constrained to the surface of the ellipsoid and moving without friction and in the absence of external forces—this path is “as straight as possible.” This allows geodesics to be extended indefinitely.
Because of their central role in geodesy, there is naturally an interest in exploring the behavior of geodesics when the earth is modeled as a triaxial ellipsoid (Burša and Fialová, 1993). In addition, there are other bodies in our solar system which are not rotationally symmetric but which can be approximated by a triaxial ellipsoid. Consequently, reliable solutions to the geodesic problems for a triaxial ellipsoid would be useful.
@@
The triaxial ellipsoid, on the other hand, possesses no obvious symmetry. It therefore came as a surprise when Jacobi (1839) found that the geodesic problem could be reduced to quadrature in this case too, with the solution given in terms of one-dimensional integrals. (We know the date, given in the abstract, for this discovery, because of a letter he wrote two days later to F. W. Bessel, his neighbor in Königsberg.) Jacobi (1843, §28) expanded on his method in his Lectures on Dynamics, and the result was generalized by Liouville (1846, §§20–21) to apply to so-called Liouville surfaces. The qualitative properties of the solution can be found in several textbooks (Darboux, 1894; Hilbert and Cohn-Vossen, 1952; Klingenberg, 1982; Arnold, 1989; Berger, 2010). However, unlike the case of the biaxial ellipsoid, where Bessel (1825) provided a prescription for computing geodesics, little effort was given to implementing Jacobi’s solution.
@@ 2 Ellipsoidal coordinates
$\mathbf{V}=\hat{\mathbf{E}}\sin\alpha+\hat{\mathbf{N}}\cos\alpha,$ (7)
where $\alpha$ is the azimuth of the geodesic measured clockwise from a line of constant $\omega$ and the hat symbol denotes a unit vector. At the pole of an oblate ellipsoid, we take the limit $\cos\beta\rightarrow 0+$, to give
@@
$\mathbf{V}=-\sin\beta(\hat{\mathbf{E}}\sin(2\alpha)-\hat{\mathbf{N}}\cos(2\alpha)),$ (12)
@@ 3 Qualitative behavior
For a biaxial ellipsoid, the equator and all the meridians are simple (not self-intersecting) closed geodesics. On the other hand, for a triaxial ellipsoid, provided it is not too eccentric, there are only three simple closed geodesics, namely the 3 principal ellipses. (For sufficiently eccentric ellipsoids, $e^{2}k^{3/2}>3/4$, other simple closed geodesics that oscillate about the major principal ellipse are possible.) The major and minor ellipses are stable; if they are perturbed, the resulting geodesic oscillates about the original ellipse. However, the median ellipse is unstable; if the geodesic is perturbed, it swings away from the $Y=0$ plane before returning to the original ellipse, but now traveling along it in the opposite direction. The stability of closed geodesics is treated in Appendix F.
@@ 4 Jacobi’s solution
$\frac{s+s_{1}}{b}$ $=\int\frac{k^{2}\cos^{2}\beta\sqrt{1-e^{2}k^{2}\cos^{2}\beta}}{\sqrt{k^{\prime 2}+k^{2}\cos^{2}\beta}\sqrt{k^{2}\cos^{2}\beta-\gamma}}\,\mathrm{d}\beta$
$\quad{}\pm\int\frac{k^{\prime 2}\sin^{2}\omega\sqrt{1+e^{2}k^{\prime 2}\sin^{2}\omega}}{\sqrt{k^{2}+k^{\prime 2}\sin^{2}\omega}\sqrt{k^{\prime 2}\sin^{2}\omega+\gamma}}\,\mathrm{d}\omega.$ (17b)
@@ 5 Properties of general geodesics
The course of a geodesic is determined by Eq. (19a) and we are immediately confronted with the problem that, for $\gamma\neq 0$, one of the integrands in Eq. (17) is singular and that, if $\gamma=0$, the integrals themselves are singular. In this section, we address the circumpolar and transpolar cases where $\gamma\neq 0$.
@@ 7 Biaxial ellipsoids
For biaxial ellipsoids, we substitute $\kappa=0$ in Eqs. (21) to give
@@
Similarly, we substitute $\kappa=1$ in Eqs. (25) to obtain
@@
In both Eqs. (42) and (43), I have replaced $\left|\mu\right|$ by $\left|\gamma\right|$ in order to facilitate their substitution into Eqs. (26). In the spherical limit, we carry out the integration to give
@@ 10 The inverse problem
  • With the biaxial problem, Helmert (1880, Eq. (6.5.1)) provided a formula for the reduced length $m_{12}$, allowing conjugate points to be found by the condition, $m_{12}=0$. I know of no corresponding formula for $m_{12}$ for the triaxial case. Nevertheless, for geodesics with $\alpha_{1}=\pm{\textstyle\frac{1}{2}}\pi$, the first conjugate point is given by the condition that $\beta$ has completed half an oscillation, which in turn implies that $\psi$ has advanced by $\pi$. Equation (26a) can be used to give the value of $\omega$ at that point. In the special case of a biaxial ellipsoid, the conjugate point for a meridional geodesic is given by Eq. (49); this is only needed for solving inverse problems on a prolate ellipsoid.
@@
11 Implementation and results
An implementation of the solutions of the direct and inverse geodesic problems is provided with version 2.7 of the C++ library, GeographicLib (Karney, 2025a). It’s practically impossible to exhaustively document the methods here; the reader is referred to the code for details. Here, I give an overview of the important aspects of the code.
@@ 11 Implementation and results
Testing for geodesics on a triaxial ellipsoid is about two orders of magnitude more challenging than the biaxial case. The shape of the ellipsoid is specified by 2 parameters ($e$ and $k$) instead of just one (the flattening), and the solution of the inverse problem depends on the longitudes of both endpoints instead of just their difference. So I limited my initial testing to a single ellipsoid, Cayley’s ellipsoid scaled to the median semiaxis, i.e., $a,b,c]=[\sqrt{2},1,1/\sqrt{2}]$, with an emphasis on exploring all the different inverse problems outlined in Sec. 10. The test set (Karney, 2024b) contains 500 000500\,000 geodesics, computed at high precision, with the coordinates of the endpoints given as integer degrees. My testing also included other ellipsoids with $a/c=2$, including oblate and prolate ellipsoids. In addition, I tested with spheres ($a=c$) with different values of $k$, and with a triaxial model of the earth (Burša and Fialová, 1993). In the following, I will only report results for the published dataset for Cayley’s ellipsoid. These should be regarded as rough indications of those likely to be obtained with ellipsoids which are not too eccentric.
@@ Appendix C Newton with bisection in two dimensions
It turns out that for the class of two-dimensional root-finding problems given by Eq. (C.1), it is quite easy to establish an axis-aligned bounding rectangle, $x\in[x_{a},x_{b}]$ and $y\in[y_{a},y_{b}]$. This starts by including all of the initial bounding parallelogram—a much looser bound. On each iteration, update one edge of the bounding rectangle depending on the signs of $f$ and $g$. Thus, in the example shown in Fig. 7, the corners of the initial bounding box are marked by circles, and the initial guess, marked by a cross, is at its center. At this point, we find that $f<0$ and $g>0$. From the figure, it is clear that the desired solution lies below this point, and therefore, we update $y_{b}$, giving a new bounding box shown as the light rectangle. This is possible because of the constraints on the slopes of the curves $f=0$ and $g=0$, which, in turn, result from the form of these functions in Eqs. (C.1). With other sign combinations for $f$ and $g$, the other edges of the bounding rectangle can be updated, and in the case where either $f$ or $g$ vanishes, two edges can be updated.
@@ Appendix E Ordinary differential equations for geodesics
Laplace (1829, §8) showed that a particle constrained to move on a surface but subject to no other forces follows a geodesic. The centrifugal acceleration of the particle is $(V^{2}/R_{c})\hat{\mathbf{U}}$, where $R_{c}$ is the radius of curvature in the direction of the velocity $\mathbf{V}$. We will take the speed $V$ to be unity (and, of course, the speed is a constant in this problem); thus, time can be replaced by $s$, the displacement along the geodesic, as the independent variable. The differential equations for the geodesic are
@@
These errors in the solution given by integrating the ordinary differential equations are an impediment to solving the inverse problem. This depends sensitively on certain properties of the solution, e.g., that a geodesic leaving an umbilic intersects the opposite umbilic. I have stitched up a solution for the inverse problem in the Octave/MATLAB version of GeographicLib by tracking the solution in ellipsoidal coordinates as the solution in cartesian coordinates unfolds (Karney, 2025b). But the code is somewhat messy, because it involves repeated coordinate conversions and has to work around the small, but inevitable, errors in the solution of the direct problem.
@@ Appendix F The stability of closed geodesics
Test data is available at Karney (2024b). Additional test data is available upon request.
@@
The C++ implementation of the algorithms given in this paper is given in GeographicLib, version 2.7, available at https://github.com/geographiclib/geographiclib/releases/tag/r2.7. This version is archived at https://doi.org/10.5281/zenodo.18011538.